OpenStudy (yanasidlinskiy):

I just need my answer checked. I have 2x+5. Let f(x) = 3x + 2 and g(x) =x - 3. Find f(x) - g(x). A. 4x - 1 B. 2x - 1 C. 2x + 5 D. 2x - 5

OpenStudy (yttrium):

correct!

OpenStudy (yanasidlinskiy):

The problem is different but it comes out to the same answer when I figured it out. And Thanx:D

OpenStudy (yanasidlinskiy):

Do you know how to solve these types of equations? Let f(x) = 3x - 6 and g(x) = x - 2. Find f/g and its domain.

OpenStudy (yanasidlinskiy):

I plugged it in and now I have \[\frac{ 3x-6 }{ x-2 }\] And now, I don't know what to do next.

OpenStudy (yttrium):

ok.

OpenStudy (yttrium):

factoring out 3 in the numerator will give you: \[\frac{ 3(x-2) }{ x-2 } = 3\]

OpenStudy (yanasidlinskiy):

It's a multiple choice question. Here's the choices: A. 3; all real numbers except x = 2 B. 3; all real numbers C. 1; all real numbers D. -3; all real numbers except x = 3

OpenStudy (yttrium):

Your target in solving for the domain of this type is that you have to make sure that your denominator will not be equal to zero (because if you make it zero, it will be undefined). When you plug in 2, it will make undefined, right? And there is no other value that will make it undefined except 2, hence, the answer is: 3; all real numbers except x = 2

OpenStudy (yanasidlinskiy):

Oh...Ok. It clearly makes sense on how you explained it:D Thank you soooooo much!!:D I wish I could go to a public shcool where someone can teach me this. But, I'm homeschooled and I'm almost done anyways. Don't have an online teacher either. I have to learn all this stuff on my own. Thanx again!!!!!:D

OpenStudy (yttrium):

Well, self-studying is also very good for it is also my means of learning. But still you need school, by the way I'm an engineering scholar. :) Okay I have to leave now. Thanks for your time :D

OpenStudy (yanasidlinskiy):

Haha, That's nice and cya:D